Mr. Thorpe inherited a sum of money. He invested 2/5 of this sum at 6%, 1/3 at 5%, and the rest at 8%. If the annual income from these investments was $2790, find the amount of money he inherited.
Found 2 solutions by VFBundy, MathTherapy:
Answer by VFBundy(438)

6% acc't:
Principal = (2/5)x = 0.40x
Rate = 0.06
Interest = 0.06(0.40x) = 0.024x

5% acc't:
Principal = (1/3)x = 0.33x
Rate = 0.05
Interest = 0.05(0.33x) = 0.0165x

8% acc't:
Principal = (1 - 2/5 - 1/3)x = (4/15)x = 0.27x
Rate = 0.08
Interest = 0.08(0.27x) = 0.0216x

0.024x + 0.0165x + 0.0216x = 2790

0.0621x = 2790

x = $44,927.54

Answer by MathTherapy(10553)


Mr. Thorpe inherited a sum of money. He invested 2/5 of this sum at 6%, 1/3 at 5%, and the rest at 8%. If the annual income from these investments was $2790, find the amount of money he inherited.
Fraction invested at 8%:  
Let amount be A
Then we get:
3(.12A) + 5(.05A) + .32A = 41,850 ------- Multiplying by LCD, 15
.36A + .25A + .32A = 41,850
.93A = 41,850
A, or amount inherited =